# Water Detection in the Interstellar Object 3I/ATLAS

### v1

@@ 3 Results and Discussion
Observations acquired with the Hubble Space Telescope placed an upper limit on the nucleus’ radius of 2.8 km (D. Jewitt et al., 2025). If the water is produced directly by sublimation of ice within the nucleus, it would require that 20% of the surface contributes, whereas most solar system comets have active fractions around 3-5% (M. F. A’Hearn et al., 1995; H. U. Keller et al., 2015).
@@
Gas detections beyond 2.5 au are uncommon due to the inefficiency of water sublimation at those distances. Only a few comets, such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), have shown OH detections at such large distances. OH emission from Bowell was detected at 5.3 au (M. F. Ahearn et al., 1984), requiring an implausibly large active area (1.6 107 km2) unless icy grains contributed significantly to water release. The OH production showed non-monotonic evolution, suggesting transient grain-driven activity. Similarly, Swift and SOHO/SWAN observations of comet Garradd revealed elevated water production beyond 3.4 au compared to narrow-slit spectroscopy, consistent with extended sources such as icy grains in the coma (M. R. Combi et al., 2013; D. Bodewits et al., 2014).
@@
Near-infrared spectroscopic observations of 3I/ATLAS obtained with Gemini South/GMOS on July 5, 2025 UTC, and IRTF/SpeX on July 14, 2025 UTC, suggest the presence of large icy grains in the coma (B. Yang et al., 2025). If these grains are the primary source of the water and OH observed in the coma, they may explain why OH (A – X) emission was detected before any CN emission. Typically, CN is among the most distant fragment species detected in comets due to the high fluorescence efficiency of its B – X (0,0) violet system near 3883 Å, which also benefits from lower atmospheric extinction and higher detector sensitivity compared to the OH emission near 3085 Å (c.f. A. L. Cochran & W. D. Cochran, 1991). Large, dark icy grains in the coma are more efficiently heated than the nucleus, leading to preferential loss of their more volatile components. A similar process was observed at 67P/Churyumov–Gerasimenko, where the transport of material from the southern to the northern hemisphere resulted in enhanced H2O abundances relative to more volatile species such as CO2, CO, and HCN, implying that these more volatile ices were lost as lofted backfall material (H. U. Keller et al., 2017).
The physical processes that govern distant cometary activity and the retention or release of volatiles in such regimes remain poorly understood. Observational biases, limited statistics, and the challenge of detecting tenuous outgassing hinder progress. However, the unusual characteristics of 3I/ATLAS including early OH but no CN detection, strong coma reddening, and inferred large grains, highlight the need for long-term, multiwavelength monitoring. Simultaneous observations of gas and ice (e.g., via JWST) will be essential to disentangle the roles of native nucleus sublimation and extended grain activity, as well as probe the abundances of any hypervolatiles like CO and CO2. As noted by both M. J. Hopkins et al. (2025) and A. G. Taylor & D. Z. Seligman (2025), 3I/ATLAS may be water-rich, consistent with planetesimals formed in the outer regions of low-metallicity planetary systems. Our results offer one of the first direct tests of that hypothesis, finding indirect evidence for a large H2O production rate, but it remains to be seen if 3I/ALTAS’ coma is dominated by the other volatiles.
The presence of substantial OH, and the inferred production of H2O, can then be used in direct comparison to 2I/Borisov. Borisov had a much lower intercept speed with our solar system and likely came from a more recently formed system; the high CO/H2O ratio post-perihelion was seen as evidence of formation near the CO ice line, likely of an M-type star (D. Bodewits et al., 2020). 3I/ATLAS, which is moving nearly twice as fast as Borisov, is also likely to show asymmetric activity around perihelion as the thermal wave penetrates into deeper reservoirs. The post-perihelion behavior of 3I will expose more pristine materials, unaltered from the inbound journey. 3I/ATLAS has a H2O production rate at 3.5 au similar to the maximum value measured for 2I/Borisov at 2.4 au (J. Crovisier et al., 2019; Z. Xing et al., 2020), and therefore we propose the following tests of the H2O-rich hypothesis:
1.
If 3I/ATLAS is H2O-rich, indicating an older age and formation in a low-metallicity system, then the H2O production rate will peak near perihelion, and only trace amounts of high-metallicity volatiles like CO,CO2, CN and hydrocarbon species will be found.
2.
If 3I/ATLAS is hyper-volatile rich (CO/CO2), indicating that its dynamic age does not necessarily represent the metallicity of where it formed, the H2O production rate will drop precipitously postperihelion as more volatile species are sublimated from depth. This would be similar to the CO/H2O observations of 2I/Borisov (D. Bodewits et al., 2020).
If the first hypothesis is born out, and 3I/ATLAS is indeed from an older low-metallicity system, than this may suggest that there is a part of the planetesimal formation or removal process that may have ejected more mass from forming systems in order to properly place the large size of 3I into context (A. G. Taylor & D. Z. Seligman, 2025). Given the ISO production of low-metallicity systems would likely come from planetary interactions (S. N. Raymond et al., 2018) or stripping by stellar flybys (S. Pfalzner et al., 2021) an overabundance of large, H2O-rich ISO’s could serve as a key indicator of ejection conditions.
@@ 4 Conclusion
We report the detection of OH (A – X) emission in the coma of 3I/ATLAS using ultraviolet imaging with the Neil Gehrels-Swift Observatory. This confirms the presence of water in the third known interstellar object. The derived production rate of molecules s-1 (40 kg s-1) is notable given the object’s large heliocentric distance of 3.51 au, where water sublimation is typically inefficient. This level of activity is consistent with a reasonably-sized active area assuming equilibrium sublimation, but could also result from the presence of large icy grains in the coma, as suggested by contemporaneous near-infrared observations. If the coma of 3I/ATLAS remains H2O dominated and depleted in higher-metallicity volatiles like CO, CN, and CO2, the relatively large size of the interstellar object may be a telling clue about the planetesimal and mass loss processes of early low-metallicity systems.
The detection of OH emission prior to any detection of CN is unusual and may indicate differences in grain composition or volatile distribution compared to solar system comets (C. Opitom et al., 2025; A. Alvarez-Candal et al., 2025). While similar behavior has been observed in comets such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), the processes governing distant activity and the release of volatiles remain poorly constrained. Grain-driven activity, in particular, may play a more significant role in interstellar objects than previously assumed. This detection demonstrates the critical role of UV space-based observations in tracking the onset and evolution of activity in distant or dynamically unusual objects. Continued monitoring of 3I/ATLAS will help clarify the thermal and compositional properties of interstellar ices and improve our understanding of volatile retention and release in planetary systems beyond our own.

### v2

@@ 4 Discussion
Using a sublimation model222https://pdssbn.astro.umd.edu/tools/ma-evap/index.shtml (J. J. Cowan & M. F. A’Hearn, 1979) and assuming constant solar elevation (as would occur for slow rotation or a Sun-facing pole), a Bond albedo of 5%, and unity infrared emissivity, we estimate the minimum required minimal active area to be 10.6 km2 for the first visit and 7.8 km2 for the second visit. Observations acquired with the Hubble Space Telescope placed an upper limit on the nucleus’s radius of 2.8 km (D. Jewitt et al., 2025). If the water is produced directly by sublimation of ice within the nucleus, it would require that 11–8% of the surface contributes, whereas most solar system comets have active fractions around 3–5% (M. F. A’Hearn et al., 1995; H. U. Keller et al., 2015).
The water production rates measured by Swift on July 31 (7.4 5.0 molecules/s) are comparable but slightly higher than the JWST measurement acquired on August 6 (2.19 0.08 molecules/s). This discrepancy may reflect systematic uncertainties if the true H2O and OH distributions deviate from the assumptions of the adopted coma models. Under such circumstances, using different aperture sizes — 3 ″for JWST and 10 ″for Swift—can yield different water production rates. For example, an extended source located outside the JWST aperture but within the Swift aperture could contribute additional OH, leading to a higher inferred production rates from the Swift measurements. The recent direct detection of water in 3I/ATLAS by JWST similarly shows a growing water production curve with nucleocentric distance, underscoring that nuclear sublimation alone cannot explain observed water activity and distributions at these heliocentric distances.
@@
Only a few comets, such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), have shown OH detections at such large distances outside of 2.5 au. OH emission from Bowell was detected at 5.3 au (M. F. A’Hearn et al., 1984), requiring an implausibly large active area (1.6 107 km2) unless icy grains contributed significantly to water release. The OH production showed non-monotonic evolution, suggesting transient grain-driven activity. Similarly, Swift and SOHO/SWAN observations of comet Garradd revealed elevated water production beyond 3.4 au compared to narrow-slit spectroscopy, consistent with extended sources such as icy grains in the coma (M. R. Combi et al., 2013; D. Bodewits et al., 2014).
@@
Near-infrared spectroscopic observations of 3I/ATLAS obtained with Gemini South/GMOS on July 5, 2025 UTC, and IRTF/SpeX on July 14, 2025 UTC, suggest the presence of large icy grains in the coma (B. Yang et al., 2025). If these grains are the primary source of the water and OH observed in the coma, they may explain why OH (A – X) emission was detected nearly simultaneously with CN emission. Typically, CN is among the most distant fragment species detected in comets due to the high fluorescence efficiency of its B – X (0,0) violet system near 3883 Å, which also benefits from lower atmospheric extinction and higher detector sensitivity compared to the OH emission near 3085 Å (c.f. A. L. Cochran & W. D. Cochran, 1991). Large, dark icy grains in the coma are more efficiently heated than the nucleus, leading to preferential loss of their more volatile components. A similar process was observed at 67P/Churyumov–Gerasimenko, where the transport of material from the southern to the northern hemisphere resulted in enhanced H2O abundances relative to more volatile species such as CO2, CO, and HCN, implying that these more volatile ices were lost as lofted backfall material (H. U. Keller et al., 2017).
With the recent spectral confirmation of CN in the coma of 3I/ATLAS we can derive log10(CN/OH) values to compare to typical solar system comet values. Reported CN production rates between August 10 and 19 span a wide range, from 0.2 to 7.81024 molecules/s (D. Schleicher, 2025; R. Rahatgaonkar et al., 2025; L. E. Salazar Manzano et al., 2025). To derive , we obtained the OH production rate () by multiplying the measured water production rate by the branching ratio (0.93) and interpolating to each , then computed for all CN measurements, yielding a median value of about , which is resilient to outliers. This would make 3I/ATLAS depleted in CN compared to even depleted solar system comets, which exhibit log10(CN/OH)=-2.69 (M. F. A’Hearn et al., 1995). However, this estimate is highly model-dependent, since is derived from water production rates using vectorial model assumptions that may not hold at large heliocentric distances.
The potential role of icy grains highlights the complexity of the physical processes governing distant activity and volatile retention or release, which remain poorly constrained due to observational biases, limited statistics, and the difficulty of detecting tenuous outgassing. Therefore, the assumptions of standard vectorial models — steady-state production, symmetric outflow, and nucleus-centered sources — may break down, leading to systematic uncertainty in the derived water and OH production rates. In addition, our preliminary radial profiles indicate a more concentrated OH distribution than predicted by the vectorial model, implying that both and may be overestimated. Consequently, the aperture-integrated number of OH molecules provides a more direct and less model-dependent constraint, while the conversion to production rates should be treated with caution.
The unusual characteristics of 3I/ATLAS including early OH, a depleted and possibly varying CN/H2O ratio (D. Schleicher, 2025; R. Rahatgaonkar et al., 2025; L. E. Salazar Manzano et al., 2025), strong coma reddening, and inferred large grains, highlight the need for long-term, multiwavelength monitoring. Simultaneous observations of gas and ice will be essential to disentangle the roles of native nucleus sublimation and extended grain activity.
@@ 5 Conclusion
We report the detection of OH (A – X) emission in the coma of 3I/ATLAS using ultraviolet imaging with the Neil Gehrels-Swift Observatory. This confirms the presence of water in the third known interstellar object. Based on a dust reddening of 29% at 4381.8Å, we obtained only a marginal detection of OH emission during the July 31–August 1 visit at au, corresponding to a water production rate of ( molecules s-1. However, subsequent observations on August 18–20 revealed a clear detection, corresponding to a water production rate of molecules s-1 (40 kg s-1) at au. This is notable given that water sublimation is typically inefficient at such large heliocentric distances. This level of activity is consistent with a reasonably-sized active area assuming equilibrium sublimation, but could also result from the presence of large icy grains in the coma, as suggested by contemporaneous near-infrared observations. If the coma of 3I/ATLAS becomes H2O-dominated during and after perihelion and shows lower abundances of higher-metallicity volatiles like CO, CN, and CO2, a trend that would be the inverse of 2I/Borisov, this may support the early and low-metallicity formation hypothesis, and the derived large size of the nucleus could be indicative of a key knowledge gap in low-metallicity system planetesimal formation and loss mechanisms.
The detection of OH emission nearly simultaneously to the detection of CN is unusual and may indicate differences in grain composition or volatile distribution compared to solar system comets (C. Opitom et al., 2025). While similar behavior has been observed in comets such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), the processes governing distant activity and the release of volatiles remain poorly constrained. Grain-driven activity, in particular, may play a more significant role in interstellar objects than previously assumed. This detection demonstrates the critical role of UV space-based observations in tracking the onset and evolution of activity in distant or dynamically unusual objects. The relatively weak CN abundances relative to H2O and the apparent large abundance of CO2 relative to H2O would seem to contradict the theorized low-metallicity formation. The upcoming perihelion passage will likely expose less-altered materials as depleted layers are stripped from its surface, and we can continue to test the low-metallicity hypothesis. Continued monitoring of 3I/ATLAS will help clarify the thermal and compositional properties of interstellar ices and improve our understanding of volatile retention and release in planetary systems beyond our own.